# Membership Inference on Text-to-image Diffusion Models via Conditional Likelihood Discrepancy

### v2

@@ 5 Related Works
Copyright protection for image-text datasets. To protect the copyright of image-text datasets, some works [44, 58, 46] incorporate imperceptible perturbations to render the released datasets unusable. Besides, other works [50, 9] utilize backdoor or watermarking to track the usage of image-text datasets. In contrast, our method demonstrates the possibility of auditing the unauthorized usage of individual image-text data points utilizing membership inference.
Membership inference on diffusion models. Recently, several membership inference methods on diffusion models have been proposed. In the grey-box or white-box setting, Carlini et al. [4] firstly conduct membership inference on unconditional diffusion models by conducting LiRA (Likelihood Ratio Attack) [3], with the requirement of training multiple shadow models. Matsumoto et al. [32] make the first step by utilizing diffusion loss to conduct query-based membership inference. Some works [11, 23] leverage the DDIM [48] deterministic forward process [22] to access the posterior estimation errors of diffusion models. And Fu et al. [13] leverage the probability fluctuations by perturbing image samples. Few works consider the black-box settings [54, 35]. However, these studies either assume partial knowledge of member set data [54] or assume extensive fine-tuning steps ($100\sim 500$ epochs) [35], both of which do not align with real-world scenarios.
@@ 6 Conclusion
In this paper, we identify the phenomenon of conditional overfitting in text-to-image models and propose CLiD-MI, the membership inference framework on text-to-image diffusion models utilizing the derived indicator, conditional likelihood discrepancy. Experimental results demonstrate the superiority of our method across various data distributions and scales, while also showing the resistance against early stopping and data augmentation strategies. Our method aims to inspire a new direction for the community regarding unauthorized usage auditing of text-image data.
@@ D.1 Monte Carlo Sampling
In Fig. D.1, we observe that the single Monte Carlo estimation achieves optimal accuracy when $t_{i}\in[400,500]$. Similar results are shown in [28]. Therefore, consistent with [28], we sample at intervals of $10$ centered around the timestep $450$. For example, if the estimation number (i.e., $M,N$ in Eq (14)) is $3$, we use the time list of $[440,450,460]$. Note that [4, 11] indicate that for DDPM of Cifar-10 [25], the best estimation timestep is around $100$. This difference may arise from the different signal-to-noise ratios of images with various resolution [18]. This finding suggests that the Monte Carlo sampling timestep should be designed differently for diffusion models of different scales.
@@ D.2 Reduction Methods
In Tab. D.1, we observe that Importance Clipping achieves the best results due to its more deterministic reduction. Additionally, we note that all three reduction methods exhibit satisfactory results, demonstrating the general applicability of our method. Comparing the results without the usage of reduction methods, the results validate the effectiveness of reduction methods in Sec. 3.5.
@@ Appendix F Ethics Statements
Appendix G Limitation
Despite the significant improvements in membership inference of text-to-image diffusion models on various data distributions and data sizes, this work still has limitations. First, due to the limited availability of open-sourced pre-trained weights of text-to-image diffusion models, evaluations under the pre-training setting are not sufficiently comprehensive. Second, the Conditional Overfitting phenomenon discussed in this paper is only explored within the domain of text-to-image diffusion models. From an insight perspective, its presence may not be such limited. Future efforts could explore this in conditional diffusion models across different modalities such as text-to-video, and further investigate more universally applicable membership inference methods.

### v3

@@ 5 Related Works
Copyright protection in text-to-image synthesis. To protect the copyright of text-to-image models, several works [68, 67] propose inserting backdoors to embed watermarks in text-to-image models. To protect the copyright of image-text datasets, some works [50, 69, 52] incorporate imperceptible perturbations to render the released datasets unusable. Other works [56, 12] utilize the backdoor or watermark to track the usage of image-text datasets. In contrast, our method indicates the possibility of auditing the unauthorized usage of individual image-text data points utilizing membership inference.
Membership inference on diffusion models. In the grey-box or white-box setting, Carlini et al. [5] firstly conduct membership inference on unconditional diffusion models by conducting LiRA (Likelihood Ratio Attack) [4], with the requirement of training multiple shadow models. Matsumoto et al. [38] make the first step by utilizing diffusion loss to conduct query-based membership inference. Some works [15, 28] leverage the DDIM [54] deterministic forward process [27] to access the posterior estimation errors of diffusion models. And Fu et al. [17] leverage the probability fluctuations by perturbing image samples. Few works consider the black-box settings [62, 41]. However, these studies either assume partial knowledge of member set data [62] or assume extensive fine-tuning steps [41] ($100\sim 500$ epochs), both of which do not align with real-world scenarios.
Memorization detection in text-to-image models. A similar work [59] detects token memorization by inspecting the magnitude of text-conditional predictions, but differs from ours by lacking in-depth rationale analysis and a rigorous membership inference setup with randomly selected member/hold-out sets.
@@ 6 Conclusion
In this paper, we identify the phenomenon of conditional overfitting in text-to-image models and propose CLiD-MI, the membership inference framework on text-to-image diffusion models utilizing the derived indicator, conditional likelihood discrepancy. Experimental results demonstrate the superiority of our method and its resistance against early stopping and data augmentation. Our method aims to inspire a new direction for the community regarding unauthorized usage auditing.
Limitations: Due to the limited availability of open-source text-to-image diffusion models, evaluations under the pretraining setting are not sufficient. Considering fine-tuning setting involves a multi step/image ratio, we acknowledge that the superiority of CLiD-MI over the baselines in the pretraining setting is not as evident compared to fine-tuning setting. We emphasize our experiments under pretraining setting (Tab. 2) reveal the hallucination success of existing works and encourage future research to focus on this more challenging and practical scenario.
Acknowledgments
We thank anonymous reviewers for their valuable feedback. In addition, we thank Xin Zhang for his editorial comments. This work is supported by the National Key R&D Program of China (No.2022YFB2703301), NSFC Projects (Nos. 92370124, 62076147). Y. Dong is also supported by the China National Postdoctoral Program for Innovative Talents and Shuimu Tsinghua Scholar Program.
@@ D.1 Monte Carlo Sampling
In Fig. D.1, we observe that the single Monte Carlo estimation achieves optimal accuracy when $t_{i}\in[400,500]$. Similar results are shown in [33]. Therefore, consistent with [33], we sample at intervals of $10$ centered around the timestep $450$. In our experiments, $M,N$ in Eq (14) are both uniformly set to $3$ (i.e., the estimation number is $3$), and we use the time list of $[440,450,460]$, resulting in the query count of 15. Note that [5, 15] indicate that for DDPM of Cifar-10 [30], the best estimation timestep is around $100$. This difference may arise from the different signal-to-noise ratios of images with various resolution [23]. This finding suggests that the Monte Carlo sampling timestep should be designed differently for diffusion models of different scales.
@@ D.2 Reduction Methods
In Tab. D.1, we observe that Importance Clipping achieves the best results due to its more deterministic reduction. So we adopt it as the reduction method used in our experiments. Additionally, we note that all three reduction methods exhibit satisfactory results, demonstrating the general applicability of our method. Comparing the results without the usage of reduction methods, the results validate the effectiveness of reduction methods in Sec. 3.5.